The average of s and t is 12, and the average of s and t and u is 15. What is the value of u?

Answer by mananth(16949)

(s+t)/2 = 12 ----------1
(s+t+u)/3 = 15 -------- 2
s+t= 12*2=24 from equation 1
substitute the vale of s+t in equation 2
(24+u) / 3 = 15
24+u = 15*3=45
u=45-24
u=21